# High Fidelity Scene Text Synthesis

### v1

@@ Abstract
Scene text synthesis involves rendering specified texts onto arbitrary images. Current methods typically formulate this task in an end-to-end manner but lack effective character-level guidance during training. Besides, their text encoders, pre-trained on a single font type, struggle to adapt to the diverse font styles encountered in practical applications. Consequently, these methods suffer from character distortion, repetition, and absence, particularly in polystylistic scenarios. To this end, this paper proposes DreamText for high-fidelity scene text synthesis. Our key idea is to reconstruct the diffusion training process, introducing more refined guidance tailored to this task, to expose and rectify the model’s attention at the character level and strengthen its learning of text regions. This transformation poses a hybrid optimization challenge, involving both discrete and continuous variables. To effectively tackle this challenge, we employ a heuristic alternate optimization strategy. Meanwhile, we jointly train the text encoder and generator to comprehensively learn and utilize the diverse font present in the training dataset. This joint training is seamlessly integrated into the alternate optimization process, fostering a synergistic relationship between learning character embedding and re-estimating character attention. Specifically, in each step, we first encode potential character-generated position information from cross-attention maps into latent character masks. These masks are then utilized to update the representation of specific characters in the current step, which, in turn, enables the generator to correct the character’s attention in the subsequent steps. Both qualitative and quantitative results demonstrate the superiority of our method to the state of the art.
@@ 1 Introduction
Scene text synthesis involves modifying or inserting specified text in arbitrary images while maintaining its natural and realistic appearance. Earlier GAN-based style transfer methods [1, 2, 3, 4] accomplish this by transferring the text style from a reference image to the rendered target text image. However, they are constrained in their capacity to generate text in arbitrary styles, such as font and color. These limitations are effectively addressed by utilizing the diffusion model [5, 6, 7, 8, 9], harnessing its inherent prior knowledge and robust capabilities. Nevertheless, these methods still encounter challenges in generating accurate text within polystylistic images due to the inadequate conditional guidance provided by their character-unaware text encoder. To this end, a recent study [10] proposes a character-level text encoder, providing more robust conditional guidance for the diffusion model.
@@
Despite their effectiveness, these methods still fail in accurately rendering text within complex scenes. Specifically, these methods such as UDiffText [10] and TextDiffuser [7] solely utilize a single font type to pre-train the text encoder, which is subsequently employed to fine-tune the generator. However, the diverse font styles encountered in practical applications present a significant challenge: the restricted representation domain hampers their ability to render text with unseen font styles, e.g. cases shown in Fig. 2 (bottom). Furthermore, these methods rely on character segmentation masks to supervise characters’ attention for position control. We argue this approach has significant limitations: (1) unlike classification tasks with uniquely determined labels, the optimal generation positions for characters can be diverse. Therefore, using the specific masks to rigidly constrain the model may limit its flexibility in estimating optimal positions, hindering its ability to adapt to varied and complex scenarios. Besides, (2) these masks generated by a pre-trained segmentation model [7] are imprecise, tending to over-segment character regions. As a result, the model encounters issues of character repetition and absence as illustrated in Fig. 2 (top left) and (top right) respectively. We suspect these issues stem from characters’ attention maps, as they reflect the underlying response to the character’s final generated position [11, 12]. Therefore, we visualize the attention maps of their problematic results by rendering attention directly onto the images. We observe certain characters’ attention misalignment, resulting in character duplication (see Fig. 3 (a) and (c)). Besides, excessive dispersion of attention also leads to character absence (see Fig. 3 (b)). This observation underscores the insufficiency of these methods in guiding the model autonomously estimating optimal character-generated position. Further, we also assess specific methods [10, 7] quantitatively on two datasets to evaluate their proficiency in directing characters’ attention toward their optimal generation regions throughout multiple global training steps. Specifically, for each test sample, we extract potential position masks for all characters based on their cross-attention maps from the inference process. Then, we compute the mean Intersection over Union (mIoU) between these masks and ground truth character segmentation masks. Their inoptimal results depicted in Fig. 4 further corroborate our hypothesis.
@@
Notably, the generator may initially struggle to direct attention to the desired generation region for each character, resulting in suboptimal latent masks and impacting training progress. Unlike existing methods [7, 10] that rigidly constrain character attention, we employ a balanced supervision strategy. Initially, we assist in calibrating the attention using character segmentation masks for a warm-up in the earlier stage. Once the model has preliminarily acquired the ability to estimate ideal generation positions, we remove this guidance, allowing the model to autonomous learning iteratively. This strategy strikes a balance between constraining the model and unleashing its flexibility in estimating optimal generation positions. Fig. 4 illustrates the efficacy of our method in continuously improving the generator’s ability to estimate the optimal characters’ position.
@@
Figure 4: The mIoU scores of UDiffText [10], TextDiffuser [7] and our method on LAION-OCR and SynthText over global training steps. Our method adopts a balanced supervision strategy: we initially use latent character masks to steer the character’s attention for a warm-up in the earlier stage and stop guiding after 25,000 steps.
Figure 5: An overview of our DreamText.
2 Related Work
Earlier GAN-based methods achieve scene text synthesis by transferring the text style in a reference image to the rendered target text image [13, 3, 14, 4]. Specifically, STEFANN [3] employs a FANnet to edit individual characters and incorporates a placement algorithm to generate the desired word. SRNet [15] and MOSTEL [1], on the other hand, split the task into two main parts: background inpainting and text style transfer. This approach allows for the end-to-end synthesizing of entire words. Despite their simplicity and effectiveness, they are limited in their ability to generate text in arbitrary styles and locations, often producing less natural-looking images.
To this end, several diffusion-based approaches have emerged to address these challenges. These approaches harness the robust capabilities of diffusion models to edit or generate scene text, thereby enhancing the quality and diversity of the generated content. DiffSTE [6] proposes a dual encoder structure, comprising a character text encoder and an instruction text encoder. These components are employed for instruction tuning, providing improved control over the backbone network. DiffUTE [16] employs an OCR-based glyph encoder to extract glyph guidance from the rendered glyph image. TextDiffuser [7] concatenates the segmentation mask as conditional input and employs character-aware loss to control the generated characters precisely. However, they still face challenges in generating accurate text within polystylistic images, primarily due to the insufficient conditional guidance provided by their character-unaware text encoder. Motivated by this, UDiffText [10] replaces the original CLIP text encoder in Stable Diffusion with a character-level text encoder with a unitary font. It also employs local attention loss for position control, relying on ground truth character segmentation maps.
@@ 2 Related Work
3 Method
Given an input image z, a text region specified by a binary mask B and a text condition c, scene text synthesis is designed to render the given text onto the input image at the given text region. A natural idea for scene text synthesis is to formulate it into a latent diffusion process [17], which is equivalent to solving the following optimization problem:
$\min_{(\theta,\vartheta)}\mathcal{L}_{LDM}\triangleq\mathbb{E}_{\textbf{{z}},c,\bm{\epsilon}\sim N(0,1),t}\parallel\bm{\epsilon}-\epsilon_{\theta}(\textbf{{z}}_{t},t,\psi_{\vartheta}(c),\textbf{{B}})\parallel_{2}^{2},$ (1)
where $\psi_{\vartheta}$ is the learnable character-level text encoder initialized by UDifftext [10] and $\theta$ is the model parameter vector. However, we argue that such a straightforward formulation could be problematic due to the following reasons:
@@ 3 Method
  • This formulation uniformly measures the mean distance between all pixels, without emphasizing the text regions. Consequently, text-unrelated regions can adversely affect the learning of the text encoder during backpropagation. Experimental analysis is provided in Sec. 4.3.
  • Due to the intricate interactions among the U-Net and text encoder, end-to-end training is particularly challenging as shown in Sec. 4.3. Therefore, a more refined guidance for the training process is necessary.
@@
3.1 Latent Diffusion with Refined Guidance
3.1.1 Latent Character Mask
@@
$\textbf{{Q}}_{l}=\textbf{{z}}_{t}\textbf{{W}}_{l}^{q},\quad\textbf{{K}}_{l}=\psi_{\vartheta}(\textit{c})\textbf{{W}}_{l}^{k},\mbox{ and }\textbf{{A}}_{l}=\mathrm{softmax}\left(\frac{\textbf{{Q}}_{l}\textbf{{K}}_{l}^{T}}{\sqrt{d}}\right),$ (2)
@@
3.1.2 Masked Diffusion Loss
We first extend the Eq. 1 to highlight all desired character tokens in the current step. Specifically, within a text c containing k tokens of interest, the diffusion loss of the corresponding pixels obtained from each token’s latent character mask is applied with an additional weighting factor $\alpha$. Formally, considering $\textbf{{M}}_{k}=\bigvee_{i=1}^{k}\textbf{{M}}_{i}$ to the union of the pixels of k characters, the masked diffusion loss [18] is given by,
$\mathcal{L}_{mask}=\mathbb{E}_{\textbf{{z}},c,\bm{\epsilon}\sim N(0,1),t}\parallel(1+\alpha\textbf{{M}}_{k})(\bm{\epsilon}-\epsilon_{\theta}(\textbf{{z}}_{t},t,\psi_{\vartheta}(c),\textbf{{B}}))\parallel_{2}^{2}.$ (5)
3.1.3 Cross Attention Loss
The masked diffusion loss is designed to focus on all desired concept tokens within a prompt. Additionally, to ensure that each token encodes information specific to its corresponding synthesis position, we incorporate the cross-attention loss [18]. This loss encourages tokens to attend exclusively to their corresponding target regions,
@@ 3.1.3 Cross Attention Loss
However, the losses mentioned above alone are inadequate for achieving optimal character representation. This is primarily due to the inherent noise in the latent character masks, which frequently results in under- or over-segmentation of the ideal character regions during training. To mitigate the influence of noise and prevent skewed learning of character representation, we introduce additional losses [10] as shown below.
3.1.4 Cross-modal Aligned Loss
@@ 3.1.5 Character Id Loss
3.2 Optimization Strategy
Heuristic Alternate Optimization: Our objective involves numerous discrete variables, rendering it non-differentiable and making vanilla SGD invalid. Therefore, we employ a heuristic alternate optimization strategy, utilizing latent character masks that encapsulate potential synthesis position information, to facilitate a symbiotic relationship between text encoder and U-Net. Specifically, we perform an alternating update between the tokens given the latent character masks and the masks themselves. During optimization, the masks are computed by Eq. 4, while for other parameters, we fix the masks and calculate gradients accordingly. Utilizing the losses above, we first optimize the representation of specific characters given the masks in each iteration, which, in turn, enables the generator to rectify the character’s attention in the subsequent steps. This iterative process enables our method to dynamically alternate between optimizing character embeddings and re-estimating character masks.
Balanced Supervision on Character Attention: Significantly, the initial training phase may challenge the generator in directing attention to the intended generation region for each character. Unlike existing methods that excessively constrain character attention which inevitably limits the model’s flexibility, we adopt a balanced supervision. Initially, we guide attention calibration by cross-entropy objective between latent character masks and character segmentation masks from datasets similar to Eq. 8 for a warm-up in the earlier stage. Once the model preliminarily acquires the capability to estimate ideal generation positions, we cease guidance, enabling it to engage in autonomous iterative learning. This strategy balances constraining the model and unleashing its flexibility in estimating characters’ optimal generation positions.
4 Experiments
4.1 Implementation Details
4.1.1 Datasets
Several datasets are employed to evaluate the competitive methods: SynthText [19], LAION-OCR [7], ICDAR13 [20] and TextSeg [21]. We provide a detailed explanation in the appendix.
4.1.2 Evaluation
Text accuracy: We assess all methods on two tasks: scene text reconstruction and scene text editing [10]. For the scene text reconstruction task, we use the models to reconstruct the text image using the provided ground truth text label and binary mask. In the scene text editing task, we replace the original text in each image with a random word of equivalent length and evaluate the models by generating images containing the edited text. We use an off-the-shelf scene text recognition (STR) model [22] to identify the rendered text and then evaluate word-level correctness using sequence accuracy (SeqAcc) by comparing the STR result with the ground truth.
Image quality: we utilize Fréchet Inception Distance (FID) [23] to quantify the distance between the text images in the dataset and the generated images. Additionally, the Learned Perceptual Image Patch Similarity (LPIPS) [24] is utilized as an additional metric.
@@ 4.1.2 Evaluation
MOSTEL [1] 0.75 0.68 0.64 0.71 0.35 0.28 0.25 0.44 25.09 0.0605
SD-Inpainting [17] 0.32 0.29 0.11 0.15 0.08 0.07 0.04 0.05 26.78 0.0696
DiffSTE [6] 0.45 0.37 0.50 0.41 0.34 0.29 0.47 0.27 51.67 0.1050
TextDiffuser [7] 0.87 0.81 0.68 0.80 0.82 0.75 0.66 0.64 32.25 0.0834
UDiffText [10] 0.94 0.91 0.93 0.90 0.84 0.83 0.84 0.78 15.79 0.0564
@@
4.2 Results
Quantitatively: In Tab. 6, we present a quantitative analysis comparing DreamText against baseline methods. Our method demonstrates a significant advantage across all quantitative metrics, highlighting the exceptional visual quality of the scene text images it generates. Notably, when compared to the previous state-of-the-art model UDiffText, DreamText outperforms it by 3.66 in terms of FID. While UDiffText and TextDiffuser excel in image quality compared to other methods, except DreamText, they tend to perform less satisfactorily in terms of sequence accuracy, especially in text editing. This discrepancy may be attributed to its restricted representation domain and deflected attention, resulting in less accurate text rendering within complex scenes.

### v2

@@ Abstract
Scene text synthesis involves rendering specified texts onto arbitrary images. Current methods typically formulate this task in an end-to-end manner but lack effective character-level guidance during training. Besides, their text encoders, pre-trained on a single font type, struggle to adapt to the diverse font styles encountered in practical applications. Consequently, these methods suffer from character distortion, repetition, and absence, particularly in polystylistic scenarios. To this end, this paper proposes DreamText for high-fidelity scene text synthesis. Our key idea is to reconstruct the diffusion training process, introducing more refined guidance tailored to this task, to expose and rectify the model’s attention at the character level and strengthen its learning of text regions. This transformation poses a hybrid optimization challenge, involving both discrete and continuous variables. To effectively tackle this challenge, we employ a heuristic alternate optimization strategy. Meanwhile, we jointly train the text encoder and generator to comprehensively learn and utilize the diverse font present in the training dataset. This joint training is seamlessly integrated into the alternate optimization process, fostering a synergistic relationship between learning character embedding and re-estimating character attention. Specifically, in each step, we first encode potential character-generated position information from cross-attention maps into latent character masks. These masks are then utilized to update the representation of specific characters in the current step, which, in turn, enables the generator to correct the character’s attention in the subsequent steps. Both qualitative and quantitative results demonstrate the superiority of our method to the state of the art. Our project page is here.
@@ Introduction
Scene text synthesis involves modifying or inserting specified text in arbitrary images while maintaining its natural and realistic appearance. Earlier GAN-based style transfer methods (Qu et al. 2023; Park et al. 2021; Roy et al. 2020; Yang, Huang, and Lin 2020) accomplish this by transferring the text style from a reference image to the rendered target text image. However, they are constrained in their capacity to generate text in arbitrary styles, such as font and color. These limitations are effectively addressed by utilizing the diffusion model (Saharia et al. 2022; Ji et al. 2023; Chen et al. 2023; Lakhanpal et al. 2024; Liu et al. 2024), harnessing its inherent prior knowledge and robust capabilities. Nevertheless, these methods still encounter challenges in generating accurate text within polystylistic images due to the inadequate conditional guidance provided by their character-unaware text encoder. To this end, a recent study (Zhao and Lian 2024) proposes a character-level text encoder, providing more robust conditional guidance for the diffusion model.
@@
Despite their effectiveness, these methods still fail in accurately rendering text within complex scenes. Specifically, these methods such as UDiffText (Zhao and Lian 2024) and TextDiffuser (Chen et al. 2023) solely utilize a single font type to pre-train the text encoder, which is subsequently employed to fine-tune the generator. However, the diverse font styles encountered in practical applications present a significant challenge: the restricted representation domain hampers their ability to render text with unseen font styles, e.g. cases shown in Fig. 2 (bottom). Furthermore, these methods rely on character segmentation masks to supervise characters’ attention for position control. We argue this approach has significant limitations: (1) unlike classification tasks with uniquely determined labels, the optimal generation positions for characters can be diverse. Therefore, using the specific masks to rigidly constrain the model may limit its flexibility in estimating optimal positions, hindering its ability to adapt to varied and complex scenarios. Besides, (2) these masks generated by a pre-trained segmentation model (Chen et al. 2023) are imprecise, tending to over-segment character regions. As a result, the model encounters issues of character repetition and absence as illustrated in Fig. 2 (top left) and (top right) respectively. We suspect these issues stem from characters’ attention maps, as they reflect the underlying response to the character’s final generated position (Wang et al. 2024a, b). Therefore, we visualize the attention maps of their problematic results by rendering attention directly onto the images. We observe certain characters’ attention misalignment, resulting in character duplication (see Fig. 3 (a) and (c)). Besides, excessive dispersion of attention also leads to character absence (see Fig. 3 (b)). This observation underscores the insufficiency of these methods in guiding the model autonomously estimating optimal character-generated position. Further, we also assess specific methods (Zhao and Lian 2024; Chen et al. 2023) quantitatively on two datasets to evaluate their proficiency in directing characters’ attention toward their optimal generation regions throughout multiple global training steps. Specifically, for each test sample, we extract potential position masks for all characters based on their cross-attention maps from the inference process. Then, we compute the mean Intersection over Union (mIoU) between these masks and ground truth character segmentation masks. Their inoptimal results depicted in Fig. 4 further corroborate our hypothesis.
@@
Notably, the generator may initially struggle to direct attention to the desired generation region for each character, resulting in suboptimal latent masks and impacting training progress. Unlike existing methods (Chen et al. 2023; Zhao and Lian 2024) that rigidly constrain character attention, we employ a balanced supervision strategy. Initially, we assist in calibrating the attention using character segmentation masks for a warm-up in the earlier stage. Once the model has preliminarily acquired the ability to estimate ideal generation positions, we remove this guidance, allowing the model to autonomous learning iteratively. This strategy strikes a balance between constraining the model and unleashing its flexibility in estimating optimal generation positions. Fig. 4 illustrates the efficacy of our method in continuously improving the generator’s ability to estimate the optimal characters’ position.
@@
Figure 4: The mIoU scores of UDiffText (Zhao and Lian 2024), TextDiffuser (Chen et al. 2023) and our method on LAION-OCR and SynthText over global training steps. Our method adopts a balanced supervision strategy: we initially use latent character masks to steer the character’s attention for a warm-up in the earlier stage and stop guiding after 25,000 steps.
Figure 5: An overview of proposed heuristic alternate optimization strategy.
Related Work
Earlier GAN-based methods achieve scene text synthesis by transferring the text style in a reference image to the rendered target text image (Kong et al. 2022; Roy et al. 2020; Shimoda et al. 2021; Yang, Huang, and Lin 2020). Specifically, STEFANN (Roy et al. 2020) employs a FANnet to edit individual characters and incorporates a placement algorithm to generate the desired word. SRNet (Wu et al. 2019) and MOSTEL (Qu et al. 2023), on the other hand, split the task into two main parts: background inpainting and text style transfer. This approach allows for the end-to-end synthesizing of entire words. Despite their simplicity and effectiveness, they are limited in their ability to generate text in arbitrary styles and locations, often producing less natural-looking images.
To this end, several diffusion-based approaches have emerged to address these challenges. These approaches harness the robust capabilities of diffusion models to edit or generate scene text, thereby enhancing the quality and diversity of the generated content. DiffSTE (Ji et al. 2023) proposes a dual encoder structure, comprising a character text encoder and an instruction text encoder. These components are employed for instruction tuning, providing improved control over the backbone network. DiffUTE (Chen et al. 2024) employs an OCR-based glyph encoder to extract glyph guidance from the rendered glyph image. TextDiffuser (Chen et al. 2023) concatenates the segmentation mask as conditional input and employs character-aware loss to control the generated characters precisely. However, they still face challenges in generating accurate text within polystylistic images, primarily due to the insufficient conditional guidance provided by their character-unaware text encoder. Motivated by this, UDiffText (Zhao and Lian 2024) replaces the original CLIP text encoder in Stable Diffusion with a character-level text encoder with a unitary font. It also employs local attention loss for position control, relying on ground truth character segmentation maps.
@@
Method
Given an input image z, a text region specified by a binary mask B and a text condition c, scene text synthesis is designed to render the given text onto the input image at the given text region. A natural idea for scene text synthesis is to formulate it into a latent diffusion process (Rombach et al. 2022), which is equivalent to solving the following optimization problem:
$\min_{(\theta,\vartheta)}\mathcal{L}_{LDM}\triangleq\mathbb{E}_{\textbf{{z}},c,\bm{\epsilon}\sim N(0,1),t}\parallel\bm{\epsilon}-\epsilon_{\theta}(\textbf{{z}}_{t},t,\psi_{\vartheta}(c),\textbf{{B}})\parallel_{2}^{2},$
where $\psi_{\vartheta}$ is the learnable character-level text encoder initialized by UDifftext (Zhao and Lian 2024) and $\theta$ is the model parameter vector. However, we argue that such a straightforward formulation could be problematic due to the following reasons:
@@
  • This formulation uniformly measures the mean distance between all pixels, without emphasizing the text regions. Consequently, text-unrelated regions can adversely affect the learning of the text encoder during backpropagation. Experimental analysis is provided in Sec. Ablation Studies and Effectiveness Analysis.
  • Due to the intricate interactions among the U-Net and text encoder, end-to-end training is particularly challenging as shown in Sec. Ablation Studies and Effectiveness Analysis. Therefore, a more refined guidance for the training process is necessary.
@@
Latent Diffusion with Refined Guidance
Latent Character Mask
@@
$\textbf{{Q}}_{l}=\textbf{{z}}_{t}\textbf{{W}}_{l}^{q},\quad\textbf{{K}}_{l}=\psi_{\vartheta}(\textit{c})\textbf{{W}}_{l}^{k},\mbox{ and }\textbf{{A}}_{l}=\mathrm{softmax}\left(\frac{\textbf{{Q}}_{l}\textbf{{K}}_{l}^{T}}{\sqrt{d}}\right),$
@@
Masked Diffusion Loss
We first extend the Eq. Method to highlight all desired character tokens in the current step. Specifically, within a text c containing k tokens of interest, the diffusion loss of the corresponding pixels obtained from each token’s latent character mask is applied with an additional weighting factor $\alpha$. Formally, considering $\textbf{{M}}_{k}=\bigvee_{i=1}^{k}\textbf{{M}}_{i}$ to the union of the pixels of k characters, the masked diffusion loss (Avrahami et al. 2023) is given by,
$\mathcal{L}_{mask}=\mathbb{E}_{\textbf{{z}},c,\bm{\epsilon}\sim N(0,1),t}\parallel(1+\alpha\textbf{{M}}_{k})(\bm{\epsilon}-\epsilon_{\theta}(\textbf{{z}}_{t},t,\psi_{\vartheta}(c),\textbf{{B}}))\parallel_{2}^{2}.$
Cross Attention Loss
The masked diffusion loss is designed to focus on all desired concept tokens within a prompt. Additionally, to ensure that each token encodes information specific to its corresponding synthesis position, we incorporate the cross-attention loss (Avrahami et al. 2023). This loss encourages tokens to attend exclusively to their corresponding target regions,
@@
However, the losses mentioned above alone are inadequate for achieving optimal character representation. This is primarily due to the inherent noise in the latent character masks, which frequently results in under- or over-segmentation of the ideal character regions during training. To mitigate the influence of noise and prevent skewed learning of character representation, we introduce additional losses (Zhao and Lian 2024) as shown below.
Cross-modal Aligned Loss
@@
CVPR’22 SD-Inpainting (Rombach et al. 2022) 0.32 0.29 0.11 0.15 0.08 0.07 0.04 0.05 26.78 0.0696
arXiv’23 DiffSTE (Ji et al. 2023) 0.45 0.37 0.50 0.41 0.34 0.29 0.47 0.27 51.67 0.1050
AAAI’23 MOSTEL (Qu et al. 2023) 0.75 0.68 0.64 0.71 0.35 0.28 0.25 0.44 25.09 0.0605
NIPS’23 TextDiffuser (Chen et al. 2023) 0.87 0.81 0.68 0.80 0.82 0.75 0.66 0.64 32.25 0.0834
ICLR’24 AnyText (Tuo et al. 2024) 0.89 0.87 0.81 0.86 0.81 0.79 0.80 0.72 22.73 0.0651
ECCV’24 UDiffText (Zhao and Lian 2024) 0.94 0.91 0.93 0.90 0.84 0.83 0.84 0.78 15.79 0.0564
@@
Optimization Strategy
Heuristic Alternate Optimization
Our objective involves numerous discrete variables, rendering it non-differentiable and making vanilla SGD invalid. Therefore, we employ a heuristic alternate optimization strategy, utilizing latent character masks that encapsulate potential synthesis position information, to facilitate a symbiotic relationship between text encoder and U-Net. Specifically, we perform an alternating update between the tokens given the latent character masks and the masks themselves. During optimization, the masks are computed by Eq. 2, while for other parameters, we fix the masks and calculate gradients accordingly. Utilizing the losses above, we first optimize the representation of specific characters given the masks in each iteration, which, in turn, enables the generator to rectify the character’s attention in the subsequent steps. This iterative process enables our method to dynamically alternate between optimizing character embeddings and re-estimating character masks.
Balanced Supervision on Character Attention
Significantly, the initial training phase may challenge the generator in directing attention to the intended generation region for each character. Unlike existing methods that excessively constrain character attention which inevitably limits the model’s flexibility, we adopt a balanced supervision. Initially, we guide attention calibration by cross-entropy objective between latent character masks and character segmentation masks from datasets similar to Eq. 5 for a warm-up in the earlier stage. Once the model preliminarily acquires the capability to estimate ideal generation positions (approximately 25k steps as shown in Fig. 4), we cease guidance, enabling it to engage in autonomous iterative learning. This strategy balances constraining the model and unleashing its flexibility in estimating characters’ optimal generation positions.
Experiments
Implementation Details
Datasets
Several datasets are employed to evaluate the competitive methods: SynthText (Gupta, Vedaldi, and Zisserman 2016), LAION-OCR (Chen et al. 2023), ICDAR13 (Karatzas et al. 2013) and TextSeg (Xu et al. 2021). We provide a detailed explanation in the appendix.
@@
Evaluation
Text accuracy: We assess all methods on two tasks: scene text reconstruction and scene text editing (Zhao and Lian 2024). For the scene text reconstruction task, we use the models to reconstruct the text image using the provided ground truth text label and binary mask. In the scene text editing task, we replace the original text in each image with a random word and evaluate the models by generating images containing the edited text. We use an off-the-shelf scene text recognition (STR) model (Bautista and Atienza 2022) to identify the rendered text and then evaluate word-level correctness using sequence accuracy (SeqAcc) by comparing the STR result with the ground truth.
Image quality: we utilize Fréchet Inception Distance (FID) (Heusel et al. 2017) to quantify the distance between the text images in the dataset and the generated images. Additionally, the Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al. 2018) is utilized as an additional metric.
Results
Quantitatively: In Tab. 6, we present a quantitative analysis comparing DreamText against baseline methods. Our method demonstrates a significant advantage across all quantitative metrics, highlighting the exceptional visual quality of the scene text images it generates. Notably, when compared to the previous state-of-the-art model UDiffText, DreamText outperforms it by 3.66 in terms of FID. While UDiffText and AnyText excel in image quality compared to other methods, except DreamText, they tend to perform less satisfactorily in terms of sequence accuracy, especially in text editing. This discrepancy may be attributed to its restricted representation domain and deflected attention, resulting in less accurate text rendering within complex scenes.